OpenStudy (anonymous):

Susi and Janet have been solving systems of equations with one polynomial function of degree two or higher and one linear function. Janet says there must always be one solutions, and Susi says there will always be two solutions. Using complete sentences, explain how Susi can be correct, how Janet can be correct, and how they both can be wrong.

OpenStudy (nurali):

because they both say always, they are both wrong. If always wasn't part of the statement, each could be correct. But if the poly has a degree of 2, and a vertex say at (0, 5) the line y = x would never intersect it. Y = x + 5 would intersect it at one point, and y = x+10 at two points.

OpenStudy (anonymous):

thank you!

OpenStudy (nurali):

Anytime.

OpenStudy (kagıtucak):

The answer depends on delta of the function. If delta greater than 0 we always have two solutions. If delta equal to 0 we always have one solution. If delta smaller than 0 we have no solution. Do you know how to find delta?

OpenStudy (anonymous):

not a clue thats in the next chapter of my flvs though :)

OpenStudy (anonymous):

but it seems that they are very important i'll put them in my notes thank you!

OpenStudy (kagıtucak):

If we have a two degree linear function like ax^2+bx+c delta=b^2-4*a*c And you can check delta to find how many solutions you have.

OpenStudy (anonymous):

ohh okay so the delta will tell me how many solutions? but how?

OpenStudy (kagıtucak):

I told it in my first comment :)

OpenStudy (anonymous):

ohh i see it but how do i find the delta of the problem?